40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jul 29 22:08:32 2010, Alberto, Update, SUS, More optics installed on the BS table 
    Reply  Fri Jul 30 02:54:04 2010, Koji, Update, 40m Upgrading, More optics installed on the BS table 
Message ID: 3326     Entry time: Thu Jul 29 22:08:32 2010     Reply to this: 3329
Author: Alberto 
Type: Update 
Category: SUS 
Subject: More optics installed on the BS table 

[Koji, Steve, Kiwamu, Alberto]

- This afternoon we installed a few new optics on the BS table: GR_PBS, GRY_SM2, GRY_SM1.

- We pulled up the cables so that we had more freedom to move one of the cable towers farther South.

- Then we re-leveled the table. PRM OSEMs were adjusted to be nominal insertions.

- Koji released the earthquake stops on BS but the readout of the OSEMs was apparently frozen on the MEDM screens.
Initially we thought it was a software problem. a nuclear reboot didn't solve it. We spent the following three hours investigating the cause.
Eventually it turned out that the earthquake stops on BS weren't actually fully released.

We opened the tank and accessed to BS. Releasing the earthquake stops in full solved the issue. The OSEMs readout went back to normal.

ELOG V3.1.3-